# Zhang, Liu, and Zhao

### v3

@@ 4.2 Convergence Analysis of Enhanced SMM Algorithm
Theorem 4.4 shows the asymptotic convergence of the enhanced SMM algorithm to composite strong d-stationarity and nonasymptotic convergence rate of the distance to the d-stationary set. We speculate that it is possible, albeit challenging, to achieve nonasymptotic results in terms of composite strong d-stationarity with carefully designed residuals and error-bound analysis, which we will leave for future study. When $\varphi$ is piecewise affine and $\Theta$ is a polytope, Assumption 2 holds naturally; furthermore, the nonasymptotic result holds without the need for Assumption 1 since a piecewise affine objective on a polytope has maximum descent rates bounded away from 0 at all points that are not d-stationary. Leveraging this special property, we can further establish the nonasymptotic convergence to d-stationarity (also local optimality in this case) of Algorithm 1 with $\varepsilon=0$, which reduces to the standard SMM with randomized index-mapping selection and achieves a faster rate with $\eta>0$. Details of the error-bound analysis and convergence proofs can be found in Appendix 10.3.
Theorem 4.5 (Convergence of SMM)
@@
For problem (9) with $\varphi$ being piecewise affine and $\Theta$ being a polytope, under the same settings of Algorithm 1 specified in Theorem 4.4 except that $\varepsilon=0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$. Moreover, with $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\mathcal{O}(T^{-\frac{1}{2}\min\left\{\alpha,1\right\}})$, which is up to a logarithmic factor when $\alpha=1$.
@@
The enhanced $\varepsilon$-active surrogation with positive $\varepsilon$ is essential for general cases. As noted in Liu et al. (2022), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity. Pang et al. (2016, Example 4) also illustrates that an MM algorithm with $\varepsilon=0$ minimizing a piecewise quadratic objective could converge to a point isolated from the d-stationary set.
@@
According to Theorems 4.4 and 4.5, the optimal iteration complexity of Algorithm 1 for $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]\leq\epsilon$ is $\mathcal{O}(\epsilon^{-2})$, which is the same as that of the stochastic DC algorithm in Le Thi et al. (2022) with incremental sampling for differentiable stochastic DC programs, provided that convex subproblems are solved to the global optimum. Notably, the optimal complexity of Algorithm 1 is achieved with the sampling rate $N_{\nu}=\mathcal{O}(\nu^{2})$ that is faster than $\mathcal{O}(\nu)$ in Le Thi et al. (2022) due to the nondifferentiable composite structure, and the convergence results of our algorithm hold for a much wider range of nonconvex and nondifferentiable composite problems. Moreover, with minor modifications of the convergence analysis, we may obtain the same convergence results for an inexact version of Algorithm 1, in which a $\delta_{\nu}$-suboptimal solution $\theta^{\nu+\frac{1}{2}}$ is computed in Step 4 and only accepted in Step 5 when it satisfies $\widehat{F}_{N_{\nu}}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{N_{\nu}}(\theta^{\nu})+\delta_{\nu}$ with $\delta_{\nu}=\mathcal{O}(\nu^{-1})$. In such a plausible extension, it is viable to design a two-loop algorithmic scheme addressing the computational complexity in solving convex subproblems. Together with our iteration complexity result, it may enable characterization of the actual work complexity of the entire algorithm similar to Pasupathy and Song (2021) and complexity comparisons with other first-order algorithms for nonconvex SP.
@@
Since the output $\widetilde{\theta}^{\,t}$ of Algorithm 1 is uniformly sampled from the iterate sequence $\{\theta^{\nu}\}_{\nu=0}^{t-1}$, the almost sure convergence of $\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,t})$ to 0 in Proposition 4.3 implies that there exists a subsequence of $\{\theta^{\nu}\}_{\nu}$ with $\bar{r}^{\varepsilon,\rho}(\theta^{\nu})$ converging to 0 almost surely. With similar convergence analysis, we can obtain the almost sure existence of the limit point of $\{\theta^{\nu}\}_{\nu}$, which satisfies the same asymptotic convergence properties as the limit points of $\{\widetilde{\theta}^{\,t}\}_{t}$.
@@
By Theorem 4.4 and the equivalence between $\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{*}$ stated in Proposition 10.2, with sufficiently large $\varepsilon$ and the proximal parameter $\eta$ close or equal to 0, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (9) with probability 1. This implies that the PADR-based ERM model together with the enhanced SMM algorithm under such settings is asymptotically optimal to the original problem (2). However, evidenced by the numerical implementations, the candidate solutions produced in Step 4 of Algorithm 1 with large $\varepsilon$ may be rejected in Step 5 very often, which results in inefficient computation and slow convergence. Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy in Section 6.
5 Extension: Constrained SP with Covariate Information
@@
6 Numerical Experiments
In this section, we evaluate the performance of the PADR-based ERM method through synthetic experiments on the Newsvendor problem and its variations with nonconvex costs or nonconvex constraints. We compare our method with PO methods and other two prominent DR-based methods. PO methods apply linear and PA regression models (referred to as PO-L and PO-PA respectively) and the squared error loss in the prediction stage; DR-based methods include NV-ERM1 method from Ban and Rudin (2018) that employs LDR and the optimizer prediction method from Bertsimas and Koduri (2021) that employs RKHS-DR. We omit comparisons with other data-driven approaches in the integrated scheme such as the objective prediction method from Bertsimas and Koduri (2021) and prescriptive methods from Bertsimas and Kallus (2019) because numerical results in Bertsimas and Koduri (2021) indicate that both two are outperformed by the optimizer prediction method in various cases.
The Newsvendor problem with covariate information $X=x$ is formulated as follows,
@@ 6 Numerical Experiments
where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$. We set the true demand as $Y=\overline{Y}(X)+\epsilon$, where $\overline{Y}(\cdot)$ is the mean model, and $\epsilon$ is a random variable with the standard normal distribution independent of $X$. Throughout all experiments, we generate the feature data uniformly from $\mathcal{X}=[-1,1]^{p}$, and we set the mean model satisfying $\min_{x\in\mathcal{X}}\overline{Y}(x)\geq 5$ to keep all corresponding demand data positive. Unless otherwise noted, both training and test data sets contain 1000 samples.
We clarify the settings throughout this section. The PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively is referred to as PADR$(K_{1},K_{2})$, and we implement multiple settings of $(K_{1},K_{2})$ from $\{0,1,2,3\}^{2}$. We set the uniform scale $\mu=50$ for parameters of PADR, PO, and LDR methods. In the implementation of the enhanced SMM algorithm, initial points are generated uniformly from $[-\mu,\mu]^{d(K_{1}+K_{2})(p+1)}$, and the sequence $\{N_{\nu}\}_{\nu}$ satisfies $N_{\nu}=\beta_{1}\nu+\beta_{2}$ with $\beta_{1},\beta_{2}>0$. Besides Algorithm 1 with constant $\varepsilon$, we also implement a version with a shrinking-$\varepsilon$ strategy, in which $\varepsilon$ is set as $\varepsilon_{0}$ at the first $T_{0}$ iterations and $\varepsilon_{1}(<\varepsilon_{0})$ at the remaining ones. We set the iterate point that obtains the minimum ERM cost in $T=10$ iterations as the output of the algorithm. For each instance, the algorithm is executed for $10$ rounds with random initialization, and we report the best performance unless otherwise noted. The hyperparameters are selected via a validation process with searching ranges listed in Table 6.
@@
Hyperparameter Selection for PADR-based ERM and Enhanced SMM. Hyperparameters Description Searching Ranges $\gamma$ the parameter for the constraints in (5) $[0,1]$ $\lambda$ the penalty parameter for exact penalization $[0,10^{3}]$ $\varepsilon_{0},\varepsilon_{1},T_{0}$ parameters for the shrinking-$\varepsilon$ strategy $[0,10^{4}],[0,10^{4}],\{1,2,\dots,6\}$ $\beta_{1},\beta_{2}$ parameters for the sampling strategy $[5,50],[10,40]$ $\eta$ the proximal parameter $[0,1]$
We use the Gaussian kernel in RKHS-DR, and the diagonal of the kernel is added by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining its performance. The kernel bandwidth, regularization parameter, and penalty parameter are selected via a validation procedure searching from $[10^{-1},10^{3}]$, $[10^{-8},1]$, and $[0,10^{3}]$ respectively. We approximate the optimal cost by the empirical cost of the theoretical (in Section 6.1) or simulated (in Sections 6.2 and 6.3) optimal decision, which is treated as the baseline and referred to as SIMOPT. For the training time comparisons, we report the total time of the enhanced SMM algorithm for 10 rounds with predetermined hyperparameters and the average time of RKHS-DR over all hyperparameter settings since the running time of RKHS-DR is highly related to the generated kernel and thus quite unstable. The training time of PO-L and LDR for each instance is less than 0.5 seconds and thus is omitted.
All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip. All programs involved are built and solved by the optimizer Gurobi, version 9.5.0.
6.1 Unconstrained Newsvendor Problem
For the Newsvendor problem (11), we set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (12)
a. Basic Results. In Figure 1a, we report comparison results with the training sample size $n$ increasing from 50 to 1000. For a fair comparison, we set the PA function class with $(K_{1},K_{2})=(3,0)$ as the regression model in PO-PA, which contains the true model $\overline{Y}(\cdot)$; the parameters of the PA regression model are estimated by solving the corresponding ERM problem via the enhanced SMM algorithm with the same hyperparameter selection procedure as PADR methods. We also implement the quadratic DR and cubic DR as the generalized LDRs, denoted by GLDR-2 and GLDR-3 respectively, the running time of which is less than 0.5 seconds and thus is omitted.
(a) Basic Results
(b) Enhanced SMM with Different $\varepsilon$-Strategies
Unconstrained Newsvendor Problem: Basic Results We also test the performance of PO-L and LDR in this instance, whose costs (around 14 and 9, respectively) are excluded in (a) for readability. In (b), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$.
@@ 6.1 Unconstrained Newsvendor Problem
The costs of PADR(3,0) and PADR(3,3) converge to SIMOPT because the optimal DR in this case is a max-affine function with three pieces, and thus PADR(3,0) and PADR(3,3) have no approximation error in principle. The RKHS-DR method converges slower than PADR(3,0) and PADR(3,3); moreover, its performance with $n=1000$ is slightly worse than that of PADR(3,0) with $n=400$. For PO-PA, given the observation that the prediction model provided by the enhanced SMM algorithm is quite close to the true model $\overline{Y}(\cdot)$, we conjecture that its gap to SIMOPT is mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$. This phenomenon is more evident from the comparison between LDR and PO-L, the two methods with misspecified models. The training time of RKHS-DR grows roughly exponentially with respect to the sample size, whereas that of PA-related methods is quite stable and significantly shorter when $n=1000$, which can be mainly credited to the sampling strategy used in the enhanced SMM algorithm. The numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted.
@@
Figure 1b displays sorted training cost curves of PADR(3,3) provided by the enhanced SMM with constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds. The shrinking-$\varepsilon$ strategy significantly improves the performance of the algorithm compared to the constant-$\varepsilon$ strategy, which typically does not reach satisfactory convergence within 10 iterations. This indicates that despite the theoretical convergence of the algorithm to global optimality with sufficiently large $\varepsilon$, it needs to engage with a huge number of index mappings when the sample size is large and thus can hardly make a valid update in each iteration. In contrast, the shrinking-$\varepsilon$ strategy with $\varepsilon_{0}=3000$ approximately attains the optimal value in 18 rounds. The significant efficiency improvement is because the shrinking-$\varepsilon$ strategy enables flexible adjustments of iterate points at early iterations as a warm start and afterward accelerates the convergence by ensuring valid updates with $\varepsilon=0$.
@@
b. Varying Feature Dimensions. We study the effect of the feature dimension $p$ on the performance of PADR methods under sparse and dense demand models $\overline{Y}$ following the settings in Bertsimas and Koduri (2021) and report the results in Figure 1d.
(c) Sparse Demand Model
(d) Dense Demand Model
Unconstrained Newsvendor Problem with Varying Feature Dimensions We also test the performance of PO-L and LDR under dense demand models, whose costs (decreasing from 15 to 5 and 9 to 4, respectively) are excluded in (b) for readability.
In the sparse case, only the first two components of $X$ contribute to $\overline{Y}$ following the demand model (12); in the dense case, we replace $X_{1}$ in (12) with the average of the first half of the features $\frac{2}{p}\sum_{i=1}^{p/2}X_{i}$ and $X_{2}$ with the average of the other half. Figure 1d implies that PADR and PO-PA are generally more robust to redundant features than RKHS-DR in the sparse case, and that PADR performs better than RKHS-DR and PO-PA in the dense case. The performance of PADR(3,3) and PADR(3,0) deteriorates as $p$ increases due to the growing estimation error and optimization difficulty. The simpler PADR(2,0) performs more stable in both cases, despite the existence of an approximation error.
c. Nonlinearity in Demand Model. We study the effect of nonlinearity in the demand model (12) with varying $k$ and report results in Figure 1f. It indicates that the performance of PADR(3,0), PADR(3,3), and PO-PA is irrelevant to the nonlinearity level of the demand model, whereas that of PADR(2,0) and RKHS-DR deteriorates as the nonlinearity level increases.
(e) $p=2\phantom{xxxxxxx}$
(f) $p=20$
Unconstrained Newsvendor Problem with Varying Nonlinearity in Demand Model In (b), the demand model is of the dense type. We also test the performance of PO-L and LDR, whose costs are excluded for readability. When $p=2$, PO-L and LDR increase from 5 to 40 and 3 to 30, respectively; when $p=20$, PO-L and LDR increase from 4 to 16 and 3 to 13, respectively.
Besides the piecewise affine demand model (12), we also test the performance of the DR-based methods with another nonlinear demand model that includes a sine function to account for seasonality. The numerical results are provided in Appendix 12.
6.2 Constrained Newsvendor Problem
We next consider a two-product Newsvendor problem with a total inventory capacity $C_{0}$. Assume that the uncensored historical demand data is available. We denote the unit back-order cost, unit holding cost, and order decision by $c_{bi}$, $c_{hi}$, and $z_{i}$ respectively for product $i=1,2$. The capacity constraint is formulated as $z_{1}+z_{2}\leq C_{0}$. In this subsection, we set $(c_{b1},c_{h1},c_{b2},c_{h2})=(8,2,2,8)$, $C_{0}=60$, and the mean demand models for the two products as
$\overline{Y}_{1}(X)=15X_{1}-5X_{2}+30,\quad\overline{Y}_{2}(X)=15X_{1}+5X_{2}+30,$
under which the probability that the constraint is active at the optimal decision is around 0.5. We compare PADR(2,2) with PO-L and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
Since DR-based methods may yield infeasible decisions, in Figure 1i we report the feasible frequency of decisions provided by PADR(2,2) and RKHS-DR. By projecting the decisions to the convex feasible region when needed, we report the test cost of the feasible decisions in Figure 1i.
(g) Basic Results
@@ 6.2 Constrained Newsvendor Problem
Newsvendor Problem with a Capacity Constraint
With the basic demand model, PADR(2,2) converges to SIMOPT and attains 100% feasibility when $n\geq 200$. In comparison, RKHS-DR and PO-L display notable gaps to SIMOPT, and the feasible frequency of RKHS-DR is only around 80%. Figures 1h and 1i display numerical results with varying feature dimensions. In the sparse case, PO-L is markedly better than the other two methods when $p=50$, indicating that regularization for sparsity is needed for learning PADR and RKHS-DR. In the dense case, PADR(2,2) performs the best except when $p=50$.
6.3 Nonconvex Newsvendor Problem
In this section, we consider two variations of the Newsvendor problem by adding a concave cost $C(z)$ in the objective or as a constraint. The cost $C(z)$ added in the objective can be interpreted as the capacity acquisition cost (Atamtürk and Hochbaum 2001); it also can be restricted in a constraint as a resource limitation. In the following experiments, we set the concave cost as
$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$
and then both two variant problems of interest are nonconvex. We compare the PADR methods with RKHS-DR and use the deterministic MM algorithm to solve the nonconvex ERM problem for learning RKHS-DR, referred to as RKHS-DR(MM). The baseline SIMOPT is computed by solving simulated nonconvex problems via a solver of mix-integer linear programming.
a. Nonconvex Objective. We first consider a generalized single-product Newsvendor with $(c_{b},c_{h})=(5,5)$ and $C(z)$ added in the objective of (11). The numerical results in Figure 1l indicate that PADR(3,0) performs the best in almost all the settings, which is consistent with the convex case in Section 6.1; moreover, the gap between PADR(3,0) and RKHS-DR(MM) in the sparse case with $p=50$ is much more significant.
(j) Basic Results
(k) Sparse Demand Model
(l) Dense Demand Model

### v4

@@ 3.2 Enhanced sampling-based MM algorithm for learning PADR
Particularly, when $\varphi$ is piecewise affine and $\Theta$ is a polytope, we develop error-bound analysis and establish the nonasymptotic convergence to d-stationarity (also local optimality in this case) of Algorithm 1 by leveraging the special property that at all points within a polytope that are not d-stationary, the piecewise affine objective has descent rates bounded away from 0. Furthermore, the nonasymptotic convergence in this case is also assured with $\varepsilon=0$, in which the algorithm reduces to the standard SMM with randomized surrogation-index selection. Details of the error-bound analysis and convergence proofs can be found in Online Appendix 8.3.
Theorem 3.7 (Nonasymptotic convergence of ESMM)
@@
For problem (10) with $\varphi$ being piecewise affine and $\Theta$ being a polytope, under the same settings of Algorithm 1 specified in Theorem 3.6 except that $\varepsilon\geq 0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$. Moreover, with $\varepsilon=0$ and $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\mathcal{O}(T^{-\frac{1}{2}\min\left\{\alpha,1\right\}})$, which is up to a logarithmic factor when $\alpha=1$.
Corollary 3.8
Under the settings specified in Theorem 3.7, to achieve $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]\leq\epsilon$, the optimal iteration complexity of Algorithm 1 is $\widetilde{\mathcal{O}}(\epsilon^{-2})$, which holds with a sampling rate $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ where $\alpha\geq 2$.
The nonasymptotic convergence to d-stationarity for general cases (including Case 2 and Case 3) can be obtained under additional assumptions on the objective $F$ and its surrogate functions, and we refer interested readers to the relative results with technical details in Online Appendix 8.4. We recognize that it is possible, albeit challenging, to achieve nonasymptotic results in terms of composite strong d-stationarity with carefully designed residuals and error-bound analysis, which we will leave for future study.
@@
The enhanced $\varepsilon$-active surrogation with positive $\varepsilon$ is essential for general cases. As noted in Liu et al. (2022), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity. This is also supported with Example 3.1 in Section 3.1.
@@
By Corollary 3.8, the iteration complexity of Algorithm 1 is similar to that of the stochastic DC algorithm in Le Thi et al. (2022) with incremental sampling for differentiable stochastic DC programs, provided that convex subproblems are solved to the global optimum. Notably, the optimal complexity of Algorithm 1 is achieved with the sampling rate that is faster than $\mathcal{O}(\nu)$ in Le Thi et al. (2022) due to the nondifferentiable composite structure, and the convergence results of our algorithm hold for a much wider range of nonconvex and nondifferentiable composite problems.
@@
With minor modifications of the convergence analysis, we may obtain the same convergence results for an inexact version of Algorithm 1, in which a $\delta_{\nu}$-suboptimal solution $\theta^{\nu+\frac{1}{2}}$ is computed in Step 4 and only accepted in Step 5 when it satisfies $\widehat{F}_{\nu}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{\nu}(\theta^{\nu})+\delta_{\nu}$ with $\delta_{\nu}=\mathcal{O}(\nu^{-1})$. In such a plausible extension, it is viable to design a two-loop algorithmic scheme addressing the computational complexity in solving convex subproblems. Together with our iteration complexity result, it may enable characterization of the actual work complexity of the entire algorithm similar to Pasupathy and Song (2021) and complexity comparisons with other first-order algorithms for nonconvex SP.
@@
Since the output $\widetilde{\theta}^{\,t}$ of Algorithm 1 is uniformly sampled from the iterate sequence $\{\theta^{\nu}\}_{\nu=0}^{t-1}$, by the almost sure convergence of $\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,t})$ to 0 in Proposition 3.5, we can easily obtain that there exists a subsequence $\{\theta^{\nu}\}_{\nu\in\kappa}$ with $\lim_{\nu(\in\kappa)\to\infty}\bar{r}^{\varepsilon,\rho}(\theta^{\nu})=0$ almost surely. With similar convergence analysis, we can obtain that the limit point of $\{\theta^{\nu}\}_{\nu\in\kappa}$ satisfies the same almost sure asymptotic convergence properties as the limit points of $\{\widetilde{\theta}^{\,t}\}_{t}$ presented in Theorem 3.6.
4 Numerical Experiments
In this section, we compare the numerical performance of the PADR-based ERM method with several benchmarks including DR-based methods, PO methods, and multiple prescriptive methods which are listed in Table 4. The experiments are conducted on Newsvendor problems with both simulated and real data and a product placement problem on a practical scale.
For the simulated data, we set the true model of $Y$ as $Y=\overline{Y}(X)+\widetilde{\epsilon}$, where $\overline{Y}(\cdot)$ represents the mean model, and $\widetilde{\epsilon}$ denotes a random variable with a standard normal distribution independent of $X$. Throughout all experiments with simulated data, we generate feature data uniformly from $[-1,1]^{p}$, and ensure the mean model satisfies $\min_{x\in[-1,1]^{p}}\overline{Y}(x)\geq 5$ which keeps all samples of $Y$ positive. Unless otherwise noted, both training and test datasets contain 1000 samples.
@@ 4 Numerical Experiments
Methods implemented in numerical experiments Abbreviation Explanation of method PADR Piecewise-affine decision rule method in this paper LDR Linear decision rule method from Ban and Rudin (2018) RKHS-DR Kernel optimizer prediction method from Bertsimas and Koduri (2021) kNN, kernel, cart, rf Prescriptive methods from Bertsimas and Kallus (2019) based on $k$-nearest neighbiors, kernels, regression trees, and random forests respectively sof Prescriptive method from Kallus and Mao (2023) based on stochastic optimization forests PO-PA, PO-L PO methods with piecewise-affine prediction model and linear prediction model respectively SIMOPT Simulated optimal decision using true conditional distribution
We refer to the PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively as PADR$(K_{1},K_{2})$. The hyperparameters $K_{1}$ and $K_{2}$ are searched within the range [0, 4] for the Newsvendor problems in Section 4.1 and Section 4.2, and [0, 6] for the product placement problem in Section 4.3. Regarding the ESMM algorithm for learning the PADR, initial points are generated uniformly from $[-50,50]^{d(K_{1}+K_{2})(p+1)}$, and the sampling sequence $\{N_{\nu}\}_{\nu}$ follows $N_{\nu}=40\nu$.The parameter $\varepsilon$ in Algorithm 1 is set to be either constant or shrinking, the latter case of which is referred to as the shrinking-$\varepsilon$ strategy. Specifically, the shrinking strategy sets the parameter a constant value $\varepsilon_{0}$ for the first $T_{0}$ iterations and then a reduced value $\varepsilon_{1}(<\varepsilon_{0})$ for the remaining ones. The ESMM algorithm produces the iterate point obtaining the minimum ERM cost as the output after $T$ iterations. For each instance, the algorithm is executed for $r$ rounds with random initialization, and we report the best performance unless otherwise noted. We set $T=10$ and $r=10$ in Sections 4.1 and 4.2, and $T=20$ and $r=5$ in Section 4.3. The reported running time of PADR is the total time of all rounds.
We clarify settings for benchmark methods listed in Table 4. We use a Gaussian kernel for kernel methods. For RKHS-DR, we adjust the diagonal of the kernel by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining performance. For PO methods, we use $\ell_{2}$ loss in the prediction stage and solve the regression problem of PO-PA using the ESMM algorithm. We set the uniform scale limit $\mu=50$ in Sections 4.1 and 4.2, and $\mu=100$ in Section 4.3. The hyperparameters of all benchmarks methods listed in Table 4 are selected via a standard validation process. The performance of LDR, PO-L, and kernel is dominated by PADR, PO-PA, and other prescriptive methods in most cases, and thus their results are omitted from the report if necessary to avoid the distraction.
@@
Hyperparameters of each method Methods Hyperparameters PADR number of pieces (PADR); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM) LDR N/A RKHS-DR kernel bandwidth, regularization parameter, penalty parameter kNN, kernel, cart, rf number of nearest neighbors (kNN); kernel bandwidth, regularization parameter (kernel); depth, minimum samples to split, minimum samples per leaf (cart and rf); sof choice of apx-soln or apx-risk method in Kallus and Mao (2023), depth, minimum samples to split, minimum samples per leaf PO-PA number of pieces (prediction model); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM solving $\ell_{2}$ regression) PO-L N/A SIMOPT N/A
All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip. All programs involved are built and solved by the optimizer Gurobi, version 9.5.0. We provide the code on GitHub: https://github.com/zhangyiy0223/PADR.
4.1 Newsvendor problem: simulated data
We consider a Newsvendor problem with covariate information $X=x$ as follows,
@@ 4.1 Newsvendor problem: simulated data
where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$. We set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (17)
4.1.1 Basic results.
@@
(d) Basic results
(e) Running time
Newsvendor problem: basic results and running time
@@
Since the optimal DR in this case is a max-affine function with three pieces, PADR$(K_{1},K_{2})$ with $K_{1}\geq 3$ has no approximation error in principle. Figure 1d shows that the costs of PADR with $(K_{1},K_{2})=(3,0)$ converge to SIMOPT as the training sample size $n$ increases from 50 to 1000. RKHS-DR converges slower than PADR; additionally, its performance with $n=1000$ is worse than that of PADR with $n=400$. Prescriptive methods, including kNN, cart, rf, and sof, are dominated by the PADR and RKHS-DR methods. For PO-PA, we observe that the prediction model provided by the PO-PA method is quite close to the true model $\overline{Y}(\cdot)$, so its substantial gap to SIMOPT could be mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$. Numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted. Figure 1e shows the training time of each method. The training time of RKHS-DR grows roughly exponentially when $n\geq 200$, whereas the time for PA-related methods remains stable and significantly shorter when $n=1000$, which is mainly credited to the enhanced and sampling techniques used in the ESMM algorithm.
[…]
(f) $\varepsilon$-Strategies
(g) Sampling technique
Performance of ESMM under various settings In (a), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$. In (b), experiments for $p=20$ are conducted on datasets using the sparse demand model introduced in Section 4.1.2. The y-axis is set on a log scale.
@@
In Figure 1g, we provide results of the ESMM algorithm under different parameter settings that illustrate the effect of the enhanced sampling strategies. Figure 1f displays curves of sorted training cost of PADR provided by the ESMM algorithm under the constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds. Under the constant-$\varepsilon$ strategy, the algorithm converges to stationary solutions which can be strengthened to global optimality with $\varepsilon$ being sufficiently large. However, Figure 1f shows that the constant-$\varepsilon$ strategy typically yields poor performance over all 30 rounds with 10 iterations. We conjecture that this is because when the sample size is large, the algorithm requires engaging with the random selection among a huge number of active indices and thus hardly makes a valid update at each iteration. In contrast, the shrinking-$\varepsilon$ strategy enables flexible adjustments of active indices for early iterations with a large positive value of $\varepsilon$ and produces iterate points serving as a warm start; afterward, it accelerates the convergence with a decreased value of $\varepsilon$. Figure 1f shows that the shrinking-$\varepsilon$ strategy has significantly better performance which attains the global optimum among roughly half of the total rounds.
Figure 1g reports the running time of PADR using the ESMM algorithm and the enhanced MM algorithm (EMM) without sampling (including all samples in the subproblem of each iteration). We use PADR(3,0) and PADR(6,4) to represent models with low and high complexity, denoted by PADRL and PADRH, respectively. The sampling technique significantly saves computational costs without compromising performance. The time consumption of ESMM is primarily determined by the sampling rate and remains relatively stable when $n\geq 400$, while the time required for EMM increases exponentially with the training sample size. The benefit of the sampling technique becomes more pronounced as the decision rule complexity increases. The time required for EMM is approximately 4 times longer than that for ESMM when $n=1000$.
4.1.2 Varying feature dimensions.
@@
We study the effect of the feature dimension of $X\in\mathbb{R}^{p}$ on the performance of PADR under sparse and dense demand models of $\overline{Y}$ following the settings in Bertsimas and Koduri (2021). Specifically, in the sparse case, only the first two components of $X$ contribute to $\overline{Y}$ following the demand model (17); in the dense case, we replace the two feature variables in (17) by the average of the first half of the features $\frac{2}{p}\sum_{i=1}^{p/2}X_{i}$ and the average of the other half respectively.
[…]
@@
Newsvendor problem with varying feature dimensions For readability, kNN is excluded under the dense demand model, whose costs exceed 8 in most trials.
Figure 1i displays the results. In the sparse case, forest-based prescriptive methods are robust to redundant features, as they only focus on informative features. When the dimension is less than 20, PADR method is competitive with comparable performance to sof and much less training time. It would be beneficial to employ an effective feature selection technique or regularization in the PADR method to improve the performance in the sparse case with a large feature dimension. In the dense case, PADR performs generally well, but its performance slightly deteriorates as $p$ increases due to growing estimation errors and optimization difficulties. We also observe that PADR with fewer pieces performs more stably in both cases despite the presence of approximation errors.
4.1.3 Nonlinearity in the demand model.
We study the effect of nonlinearity of the demand with respect to the feature variables by varying the coefficient $k$ in the demand model (17) and report results in Figure 1k. It indicates that the performance of PADR and PO-PA is irrelevant to the nonlinearity level of the demand model due to the strong approximation ability to the nonlinear structure. In contrast, the performance of RKHS-DR and prescriptive methods deteriorates significantly as the nonlinearity level increases. These methods require significantly more samples to achieve the similar performance when the level of nonlinearity increases.
[…]
(j) $p=2\phantom{xxxxxxx}$
(k) $p=20$
Newsvendor problem with varying nonlinearity in demand model In (b), the demand model is of the dense type. For readability, kNN, with costs increasing from 5 to 40, is excluded in (a).